OpenStudy (anonymous):

sue bought a pair of jeans and a shirt for $55.75 if the jeans cost $25.25 how much did she pay for the shirt

OpenStudy (jack1):

eqn 1: j + s = 55.75 eqn 2: j = 25.25 so sub eqn 2 into eqn 1

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I GOT 25.25 IS THAT CORRECT

OpenStudy (jack1):

no...

OpenStudy (jack1):

j + s = 55.75 25.25 + s = 55.75 so s = 55.75 - 25.25 =...?

OpenStudy (anonymous):

I GOT 30.50

OpenStudy (jack1):

sweet, spot on!

OpenStudy (anonymous):

OK WHEN 5 IS ADDED TO 2 TIMES A NUMBER THE RESULT IS 45 I NEED HELP WITH THIS QUETION

OpenStudy (jack1):

2x + 5 = 45 ...solve for x

OpenStudy (anonymous):

I GOT 20

OpenStudy (jack1):

yep, perfect

OpenStudy (jack1):

and turn off caps lock, it seems like you're always yelling ;P

OpenStudy (anonymous):

i need help with lats quetion in a given amount of time, mary drove twice as far as lynn. altogether they drove 60 miles how far did mary drive